find two consecutive integers whose sum is 83
Answer by xcentaur(357)

Let the first integer be x.
Second integer = (x+1)
"Consecutive integers sum up to 83"
This statement gives us,
x+(x+1)=83
2x+1=83
2x=82
x=41


Then first number is 41,and second number is 42.
41+42=83 ... therefore these are the required integers.


ANS: 41,42


Hope this helps,
best of luck.